karlieeeeeee:

Is (10, –6) a solution to this system of equations? y= 1/5x-8 y= -1/2x-1

Angle:

Hello @karlieeeeeee ! Welcome to QuestionCove For a problem like this, the main strategy is to plug in the values of (10, -6) into each equation and see what you get so plug in ` x = 10 ` and ` y = -6 ` into ` y = 1/5 x - 8 `

Angle:

In this case, we would get (-6) = (1/5)(10) - 8 The question becomes: "Does (1/5)(10) - 8 equal -6?"

Angle:

Next we do the same thing with the other equation plug in ` x = 10 ` and ` y = -6 ` into ` y = -1/2 x - 1` then check if the left side is equal to the right side

Angle:

if both results are true, then your answer would be "yes, (-10, -6) is a solution) if one of the results is false, then your answer would be "no"

Angle:

@karlieeeeeee if you come back and are still stuck, just let me know and we can take this one step at a time

karlieeeeeee:

So would it be yes or no?

Angle:

"Does (1/5)(10) - 8 equal -6?"

karlieeeeeee:

yes that's the question

Angle:

nono, I'm asking YOU what is (1/5)(10) - 8 equal to?

karlieeeeeee:

I don't know that's why I'm asking for help

Angle:

What is (1/5) times 10 equal to?

Angle:

\(\frac{1}{5}\times10\) = ?

karlieeeeeee:

2

Angle:

What is 2 minus 8?

karlieeeeeee:

-6

Angle:

is -6 equal to -6?

karlieeeeeee:

yes

Angle:

Awesome, this means that ` (10, -6) ` is a possible solution for ` y = (1/5)x - 8 ` now we check for ` y = (-1/2)x - 1 ` What is \(\frac{-1}{2}(10)-1\) = ?

karlieeeeeee:

-6

Angle:

is -6 = -6 ?

karlieeeeeee:

yes

Angle:

ok, so because we answered "yes" for BOTH equations your answer is going to be yes for your problem

karlieeeeeee:

may you please help me with more?

Angle:

We prefer to keep it to one question per question thread. You are welcome to click the "close" button on this question and open a new question post

karlieeeeeee:

so if I post another question you will help?